### v1

# A note on Combinatorial Invariance of Kazhdan–Lusztig polynomials

Abstract.
We introduce the concepts of an amazing hypercube decomposition and a double shortcut for it, and use these new ideas to formulate a conjecture implying the Combinatorial Invariance Conjecture of the Kazhdan–Lusztig polynomials for the symmetric group. This conjecture has the advantage of being combinatorial in nature.
Key words and phrases:
Kazhdan-Lusztig polynomials, Combinatorial invariance, hypercube decompositions
2020 Mathematics Subject Classification:
05E10 - 05E16 (primary), 20F55 (secondary)
1. Introduction
The Combinatorial Invariance Conjecture for Kazhdan–Lusztig polynomials was formulated independently by Lusztig and Dyer in the ’80’s. Its version for the symmetric group $W$ is the following.
Conjecture 1.1.
Let $u,v\in W$. The Kazhdan–Lusztig polynomial $P_{u,v}(q)$ (or, equivalently, the Kazhdan–Lusztig $\widetilde{R}$-polynomial $\widetilde{R}_{u,v}(q)$) depends only on the isomorphism class of the Bruhat interval $[u,v]$ as a poset.
It has been the focus of active research for the past forty years and has recently received important new inputs by [3] and [6], where the concept of a hypercube decomposition was introduced and used to give a conjectural formula for the Kazhdan–Lusztig polynomials of the symmetric group that implies Conjecture 1.1.
Successively, a conjectural formula for $\widetilde{R}$-polynomial $\widetilde{R}_{u,v}(q)$ of the symmetric group was given in terms of shortcuts ([5]) of a hypercube decomposition. Also this formula would imply Conjecture 1.1. A further work studying hypercube decompositions is [1], where the authors give another conjecture implying Conjecture 1.1 and prove combinatorial invariance for elementary intervals, which marks an improvement of the result in [4] for lower Bruhat intervals.
In this short note, we introduce the concepts of an amazing hypercube decomposition and a double shortcut for it, and use these new ideas to formulate yet another conjecture implying the conjecture in [5] and thus ultimately Conjecture 1.1. This new conjecture has the advantage of being combinatorial in nature, with no mention of Kazhdan–Lusztig polynomials: a purely graph theoretic property of the Bruhat interval would imply the Combinatorial Invariance Conjecture for the symmetric group.
The conjecture presented in this paper has been verified (in a strong form) up to $S_{6}$ by computer calculations.
2. Preliminaries
In this section we recall the basic definitions and set the notation to be used. We refer the reader to [2] for undefined terminology concerning Coxeter groups.
We recall the definition of a hypercube decomposition from [3]. Let $(W,S)$ be a Coxeter group of type $A$. We denote by $B(W)$ the Bruhat graph of $W$.
Given a set $E$, let ${\mathcal{P}}(E)$ denote the directed Boolean algebra on $E$, i.e. the directed graph having the power set of $E$ as vertex set and where $I\rightarrow J$ if $I$ is obtained from $J$ by removing one element.
Let $p\in W$ and $E$ be a set of arrows of $B(W)$ all having $p$ as target. Then $E$ spans a hypercube if there exists a unique embedding of directed graphs $\theta:{\mathcal{P}}(E)\rightarrow W$ sending the directed edge $E\setminus{\{\alpha\}}\rightarrow E$ to $\alpha$, for all $\alpha\in E$. Furthermore, $E$ spans a hypercube cluster if every subset $E^{\prime}$ of $E$ consisting of edges with pairwise incomparable sources (with respect to Bruhat order) spans a hypercube.
Let $u,v\in W$ and $z\in[u,v]$. We say that $z$ is an upper hypercube decomposition of $[u,v]$ provided that:
  1. (1)
    $[z,v]$ is diamond complete (with respect to $[u,v]$), meaning that, if there exist $x\in[u,v]$ and $a_{1},a_{2},y\in[z,v]$, $a_{1}\neq a_{2}$, such that $x\rightarrow a_{1}\rightarrow y$, $x\rightarrow a_{2}\rightarrow y$, then $x\in[z,v]$;
  2. (2)
    for all $p\in[z,v]$, the set $E^{p}=\{x\rightarrow p:x\notin[z,v]\}$ spans a hypercube cluster.
Similarly, taking the dual versions of the above definitions, we obtain the concept of a lower hypercube decomposition.
Following [5], we give the next definition. We denote by $\operatorname{d}(x,y)$ the distance function of $B(W)$, i.e., the minimum of the lengths of paths from $x$ to $y$.
Definition 2.1.
Let $u,v,z\in W$ with $z\in[u,v]$. We let
$W_{[u,v]}^{z}:=\{p\in[z,v]:\operatorname{supp}(\Gamma)\cap[z,v]=\{p\}\textrm{ for all paths $\Gamma$ from $u$ to $p$ of length $\operatorname{d}(u,p)$}\},$
and
$\widetilde{R}_{u,v}^{z}(q):=\sum_{p\in W_{[u,v]}^{z}}q^{\operatorname{d}(u,p)}\widetilde{R}_{p,v}(q).$
We call the elements in $W_{[u,v]}^{z}$ the (upper) shortcuts of $[u,v]$ with respect to $z$. Furthermore, we say that an element $z$ in $[u,v]$ is an (upper) $R$-element for $[u,v]$ if $\widetilde{R}_{u,v}^{z}=\widetilde{R}_{u,v}$.
Remark 2.2.
The elements $\min([u,v]\cap W_{S\setminus\{s_{n-1}\}}v)$, $\min([u,v]\cap W_{S\setminus\{s_{1}\}}v)$, $\min([u,v]\cap vW_{S\setminus\{s_{n-1}\}})$, and $\min([u,v]\cap vW_{S\setminus\{s_{1}\}}v)$ are always upper hypercube decompositions for the interval $[u,v]$ (possibly not mutually distinct). We call these elements the standard upper hypercube decompositions for $[u,v]$. Note that these hypercube decompositions are called canonical in [5].
Remark 2.3.
We note that the concept of an upper shortcut can be equivalently defined in the following way. Let $u,v\in W$, and $z$ be an upper hypercube decomposition of $[u,v]$. An element $p$ in $[z,v]$ is an upper shortcut of $[u,v]$ with respect to $z$ if $\operatorname{d}(u,p)<\operatorname{d}(u,x)$ for all $x$ in $[z,p]$ satisfying $\operatorname{d}(x,p)=1$.
3. Amazing hypercube decompositions and double shortcuts
Definition 3.1.
Let $u,v\in W$. We say that an upper hypercube decomposition $z$ of $[u,v]$ is amazing provided that, for all $x\in[u,v]$, the intersection $[z,v]\cap[x,v]$ has a minimum $z\wedge x$, and this minimum $z\wedge x$ is a (necessarily amazing) hypercube decomposition of $[x,v]$. Furthermore, we say that an amazing hypercube decomposition $z$ of $[u,v]$ is an (upper) amazing $R$-element for $[u,v]$ if $z\wedge x$ is an (upper) $R$-element for $[x,v]$ for all $x\in[u,v]$.
Remark 3.2.
Let $u,v\in W$ and $z$ be a standard hypercube decomposition of $[u,v]$. Since $z\wedge x$ exists and is a standard hypercube decomposition of $[x,v]$ (of the same kind), we have that $z$ is an amazing hypercube decomposition, and, by [5, Corollary 3.10], an amazing $R$-element.
The following conjecture is a weak version of Conjecture of [5], which has been verified up to $W=S_{6}$.
Conjecture 3.3.
Let $W$ be a Coxeter group of type $A$, $u,v\in W$, and $z\in[u,v]$ be an amazing upper hypercube decomposition. Then $z$ is an $R$-element, i.e.
$\widetilde{R}_{u,v}=\sum_{p\in W_{[u,v]}^{z}}q^{\operatorname{d}(u,p)}\widetilde{R}_{p,v}.$
Conjecture 3.3 implies the Combinatorial Invariance Conjecture.
Definition 3.4.
Let $u,v\in W$, and $z$ and $z^{\prime}$ be two amazing upper hypercube decomposition of $[u,v]$. We say that an element $b$ in $[z,v]\cap[z^{\prime},v]$ is a $(z,z^{\prime})$-double shortcut of $[u,v]$ if there exists $p\in W^{z}_{[u,v]}$ such that $b\in W^{z\wedge z^{\prime}}_{[p,v]}$. We denote by $DS(z,z^{\prime})$ the multiset
$\{(\operatorname{d}(u,p)+\operatorname{d}(p,b),b):p\in W^{z}_{[u,v]}\text{ and}b\in W^{z\wedge z^{\prime}}_{[p,v]}\}.$
Theorem 3.5.
Let $W$ be a Coxeter group of type $A$, $u,v\in W$, and $z,z^{\prime}\in[u,v]$ be two amazing upper hypercube decomposition. Suppose that
  1. (1)
    $z$ is an amazing $R$-element,
  2. (2)
    $z^{\prime}\wedge x$ is an (amazing) $R$-element for all $x\in[u,v]\setminus\{u\}$,
  3. (3)
    $DS(z,z^{\prime})=DS(z^{\prime},z)$ (as multisets).
Then, $z^{\prime}$ is an (amazing) $R$-element.
Proof.
We have:
$\widetilde{R}_{u,v}$ $=$ $\sum_{p\in W_{[u,v]}^{z}}q^{\operatorname{d}(u,p)}\widetilde{R}_{p,v}$
$=$ $\sum_{p\in W_{[u,v]}^{z}}q^{\operatorname{d}(u,p)}\big{(}\sum_{b\in W_{[p,v]}^{z^{\prime}\wedge p}}q^{\operatorname{d}(p,b)}\widetilde{R}_{b,v}\big{)}$
$=$ $\sum_{(a,b)\in DS(z,z^{\prime})}q^{a}\widetilde{R}_{b,v}$
$=$ $\sum_{(a,b)\in DS(z^{\prime},z)}q^{a}\widetilde{R}_{b,v}$
$=$ $\sum_{p^{\prime}\in W_{[u,v]}^{z^{\prime}}}q^{\operatorname{d}(u,p^{\prime})}\big{(}\sum_{b\in W_{[p^{\prime},v]}^{z\wedge p^{\prime}}}q^{\operatorname{d}(p^{\prime},b)}\widetilde{R}_{b,v}\big{)}$
$=$ $\sum_{p^{\prime}\in W_{[u,v]}^{z^{\prime}}}q^{\operatorname{d}(u,p^{\prime})}\widetilde{R}_{p^{\prime},v},$
where the first equation follows by (1), the second equation follows by (2), the third equation by the definition of the multiset $DS(z,z^{\prime})$, the fourth by (3), the fifth by the definition of the multiset $DS(z^{\prime},z)$, and the sixth equation follows by (1). ∎
Fix $u,v\in W$. Let us define an equivalence relation on the set of amazing hypercube decompositions of $[u,v]$ as the transitive closure of the relation $z\sim z^{\prime}$ if $DS(z,z^{\prime})=DS(z^{\prime},z)$.
Conjecture 3.6.
Let $W$ be a Coxeter group of type $A$. The above relation is trivial (i.e., it has only one equivalence class).
The following conjecture is a weakening of Conjecture 3.6.
Conjecture 3.7.
Let $W$ be a Coxeter group of type $A$. Every equivalence class of the above relation contains an amazing $R$-element.
Remark 3.8.
Up to $S_{6}$, we checked by computer that, in fact, a much stronger statement than Conjecture 3.7 holds: $DS(z,z^{\prime})=DS(z^{\prime},z)$ for all amazing hypercube decompositions of any interval.
Theorem 3.9.
Conjecture 3.7 implies Conjecture 3.3.
Proof.
Towards a contradiction, let $[u,v]$ be an interval of minimal length having an amazing hypercube decomposition $z$ thath fails to be an $R$-element. By hypothesis, there exists a sequence $(z_{0},\ldots,z_{r})$ of amazing hypercube decompositions such that $z_{0}=z$, and $z_{r}$ is an amazing $R$-element, and $DS(z_{i-1},z_{i})=DS(z_{i},z_{i-1})$ for each $i\in[r]$. One gets a contradiction by iteratively applying Theorem 3.5. ∎
Acknowledgments: We would like to thank Grant T. Barkley, Francesco Brenti, and Christian Gaetz for stimulating discussions we had during the Workshop “Bruhat order: recent developments and open problems”held at the University of Bologna from 15 April 2024 to 19 April 2024. The authors are members of the I.N.D.A.M. group GNSAGA.

### v3

@@
 This conjecture has the advantage of being combinatorial in nature.
+The appendix by Grant T.
+Barkley and Christian Gaetz discusses the related notion of double hypercubes and proves an analogous conjecture for these in the case of co-elementary intervals.
 Key words and phrases:
@@
 The Kazhdan–Lusztig polynomial $P_{u,v}(q)$ (or, equivalently, the Kazhdan–Lusztig $\widetilde{R}$-polynomial $\widetilde{R}_{u,v}(q)$) depends only on the isomorphism class of the Bruhat interval $[u,v]$ as a poset.
-It has been the focus of active research for the past forty years and has recently received important new inputs by [3] and [6], where the concept of a hypercube decomposition was introduced and used to give a conjectural formula for the Kazhdan–Lusztig polynomials of the symmetric group that implies Conjecture 1.1.
+It has been the focus of active research for the past forty years (we refer the reader to [6] for more details) and has recently received important new inputs by [3] and [7], where the concept of a hypercube decomposition was introduced and used to give a conjectural formula for the Kazhdan–Lusztig polynomials of the symmetric group that implies Conjecture 1.1.
 Successively, a conjectural formula for $\widetilde{R}$-polynomial $\widetilde{R}_{u,v}(q)$ of the symmetric group was given in terms of shortcuts ([5]) of a hypercube decomposition.
@@
 A further work studying hypercube decompositions is [1], where the authors give another conjecture implying Conjecture 1.1 and prove combinatorial invariance for elementary intervals, which marks an improvement of the result in [4] for lower Bruhat intervals.
-In this short note, we introduce the concepts of an amazing hypercube decomposition and a double shortcut for it, and use these new ideas to formulate yet another conjecture implying the conjecture in [5] and thus ultimately Conjecture 1.1.
+A new different approach (via flipclasses) is proposed in [9].
+In this paper, we introduce the concepts of an amazing hypercube decomposition and a double shortcut for it, and use these new ideas to formulate yet another conjecture implying the conjecture in [5] and thus ultimately Conjecture 1.1.
 This new conjecture has the advantage of being combinatorial in nature, with no mention of Kazhdan–Lusztig polynomials: a purely graph theoretic property of the Bruhat interval would imply the Combinatorial Invariance Conjecture for the symmetric group.
@@
 In this section we recall the basic definitions and set the notation to be used. We refer the reader to [2] for undefined terminology concerning Coxeter groups.
-We recall the definition of a hypercube decomposition from [3]. Let $(W,S)$ be a Coxeter group of type $A$. We denote by $B(W)$ the Bruhat graph of $W$.
+We recall the definition of a hypercube decomposition from [3]. Let $(W,S)$ be a Coxeter group of type $A_{n}$, i.e., the symmetric group $S_{n+1}$. We denote by $B(W)$ the Bruhat graph of $W$.
 Given a set $E$, let ${\mathcal{P}}(E)$ denote the directed Boolean algebra on $E$, i.e.
@@
 Similarly, taking the dual versions of the above definitions, we obtain the concept of a lower hypercube decomposition.
+We introduce the $\widetilde{R}$-polynomials by the following combinatorial interpretation due to Dyer (see [8] and also [2, Theorem 5.3.4]).
+Given a path $\Gamma=(x_{0}\rightarrow x_{1}\rightarrow\cdots\rightarrow x_{r})$, we denote by $|\Gamma|$ the length of $\Gamma$ (i.e., $|\Gamma|=r$), and we define $\{x_{i}:i\in[0,r]\}$ to be the support of $\Gamma$, denoted $\operatorname{supp}(\Gamma)$.
+Theorem 2.1.
+Let $\preceq$ be a reflection ordering, and $u,v\in\mathfrak{S}_{n}$. Then
+$\widetilde{R}_{u,v}(q)=\sum q^{|\Gamma|},$
+where the sum is over all increasing paths $\Gamma$ from $u$ to $v$.
 Following [5], we give the next definition. We denote by $\operatorname{d}(x,y)$ the distance function of $B(W)$, i.e., the minimum of the lengths of paths from $x$ to $y$.
@@
 Let $u,v,z\in W$ with $z\in[u,v]$. We let
-$W_{[u,v]}^{z}:=\{p\in[z,v]:\operatorname{supp}(\Gamma)\cap[z,v]=\{p\}\textrm{ for all paths $\Gamma$ from $u$ to $p$ of length $\operatorname{d}(u,p)$}\},$
+$W_{[u,v]}^{z}:=\{p\in[z,v]:\operatorname{supp}(\Gamma)\cap[z,v]=\{p\}\textrm{ for all paths $\Gamma$ from $u$ to $p$ with $|\Gamma|=\operatorname{d}(u,p)$}\},$
 and
@@
 3. Amazing hypercube decompositions and double shortcuts
+The following definition strengthens the definition of a join hypercube decomposition appearing in [5] (indeed, it is a stable version of it).
 Definition 3.1.
 Let $u,v\in W$.
-We say that an upper hypercube decomposition $z$ of $[u,v]$ is amazing provided that, for all $x\in[u,v]$, the intersection $[z,v]\cap[x,v]$ has a minimum $z\wedge x$, and this minimum $z\wedge x$ is a (necessarily amazing) hypercube decomposition of $[x,v]$.
-Furthermore, we say that an amazing hypercube decomposition $z$ of $[u,v]$ is an (upper) amazing $R$-element for $[u,v]$ if $z\wedge x$ is an (upper) $R$-element for $[x,v]$ for all $x\in[u,v]$.
+We say that an upper hypercube decomposition $z$ of $[u,v]$ is amazing provided that, for all $x\in[u,v]$, the intersection $[z,v]\cap[x,v]$ has a minimum $z\vee x$, and this minimum $z\vee x$ is a (necessarily amazing) hypercube decomposition of $[x,v]$.
+Furthermore, we say that an amazing hypercube decomposition $z$ of $[u,v]$ is an (upper) amazing $R$-element for $[u,v]$ if $z\vee x$ is an (upper) $R$-element for $[x,v]$ for all $x\in[u,v]$.
 Remark 3.2.
 Let $u,v\in W$ and $z$ be a standard hypercube decomposition of $[u,v]$.
-Since $z\wedge x$ exists and is a standard hypercube decomposition of $[x,v]$ (of the same kind), we have that $z$ is an amazing hypercube decomposition, and, by [5, Corollary 3.10], an amazing $R$-element.
-The following conjecture is a weak version of Conjecture of [5], which has been verified up to $W=S_{6}$.
+Since $z\vee x$ exists and is a standard hypercube decomposition of $[x,v]$ (of the same kind), we have that $z$ is an amazing hypercube decomposition, and, by [5, Corollary 3.10], an amazing $R$-element.
+The following conjecture is a weak version of Conjecture 6.1 of [5], which has been verified up to $W=S_{6}$.
 Conjecture 3.3.
@@
 Let $u,v\in W$, and $z$ and $z^{\prime}$ be two amazing upper hypercube decomposition of $[u,v]$.
-We say that an element $b$ in $[z,v]\cap[z^{\prime},v]$ is a $(z,z^{\prime})$-double shortcut of $[u,v]$ if there exists $p\in W^{z}_{[u,v]}$ such that $b\in W^{z\wedge z^{\prime}}_{[p,v]}$.
+We say that an element $b$ in $[z,v]\cap[z^{\prime},v]$ is an (upper) $(z,z^{\prime})$-double shortcut of $[u,v]$ if there exists $p\in W^{z}_{[u,v]}$ such that $b\in W^{z^{\prime}\vee p}_{[p,v]}$.
 We denote by $DS(z,z^{\prime})$ the multiset
-$\{(\operatorname{d}(u,p)+\operatorname{d}(p,b),b):p\in W^{z}_{[u,v]}\text{ and}b\in W^{z\wedge z^{\prime}}_{[p,v]}\}.$
+$\{(\operatorname{d}(u,p)+\operatorname{d}(p,b),b):p\in W^{z}_{[u,v]}\text{ and}b\in W^{z^{\prime}\vee p}_{[p,v]}\}.$
+Figure 1 depicts an interval $[u,v]$, two amazing upper hypercube decompositions $z$ and $z^{\prime}$ of $[u,v]$, an upper shortcut $p$ of $[u,v]$ with respect to $z$, and an upper shortcut $b$ of $[p,v]$ with respect to $z^{\prime}\vee p$.
+The element $b$ is hence a $(z,z^{\prime})$-double shortcut of $[u,v]$.
+[…]
+Figure 1. A $(z,z^{\prime})$-double shortcut of $[u,v]$.
 Theorem 3.5.
@@
   2. (2)
-    $z^{\prime}\wedge x$ is an (amazing) $R$-element for all $x\in[u,v]\setminus\{u\}$,
+    $z^{\prime}\vee x$ is an (amazing) $R$-element for all $x\in[u,v]\setminus\{u\}$,
   3. (3)
@@
 $\widetilde{R}_{u,v}$ $=$ $\sum_{p\in W_{[u,v]}^{z}}q^{\operatorname{d}(u,p)}\widetilde{R}_{p,v}$
-$=$ $\sum_{p\in W_{[u,v]}^{z}}q^{\operatorname{d}(u,p)}\big{(}\sum_{b\in W_{[p,v]}^{z^{\prime}\wedge p}}q^{\operatorname{d}(p,b)}\widetilde{R}_{b,v}\big{)}$
+$=$ $\sum_{p\in W_{[u,v]}^{z}}q^{\operatorname{d}(u,p)}\big{(}\sum_{b\in W_{[p,v]}^{z^{\prime}\vee p}}q^{\operatorname{d}(p,b)}\widetilde{R}_{b,v}\big{)}$
 $=$ $\sum_{(a,b)\in DS(z,z^{\prime})}q^{a}\widetilde{R}_{b,v}$
 $=$ $\sum_{(a,b)\in DS(z^{\prime},z)}q^{a}\widetilde{R}_{b,v}$
-$=$ $\sum_{p^{\prime}\in W_{[u,v]}^{z^{\prime}}}q^{\operatorname{d}(u,p^{\prime})}\big{(}\sum_{b\in W_{[p^{\prime},v]}^{z\wedge p^{\prime}}}q^{\operatorname{d}(p^{\prime},b)}\widetilde{R}_{b,v}\big{)}$
+$=$ $\sum_{p^{\prime}\in W_{[u,v]}^{z^{\prime}}}q^{\operatorname{d}(u,p^{\prime})}\big{(}\sum_{b\in W_{[p^{\prime},v]}^{z\vee p^{\prime}}}q^{\operatorname{d}(p^{\prime},b)}\widetilde{R}_{b,v}\big{)}$
 $=$ $\sum_{p^{\prime}\in W_{[u,v]}^{z^{\prime}}}q^{\operatorname{d}(u,p^{\prime})}\widetilde{R}_{p^{\prime},v},$
@@
 Proof.
-Towards a contradiction, let $[u,v]$ be an interval of minimal length having an amazing hypercube decomposition $z$ thath fails to be an $R$-element.
+Towards a contradiction, let $[u,v]$ be an interval of minimal length having an amazing hypercube decomposition $z$ that fails to be an $R$-element.
 By hypothesis, there exists a sequence $(z_{0},\ldots,z_{r})$ of amazing hypercube decompositions such that $z_{0}=z$, and $z_{r}$ is an amazing $R$-element, and $DS(z_{i-1},z_{i})=DS(z_{i},z_{i-1})$ for each $i\in[r]$.
@@
 ∎
+The following result supports Conjecture 3.7.
+Theorem 3.10.
+Let $(W,S)$, $(W_{1},S_{1})$, and $(W_{2},S_{2})$ be three Coxeter systems of type $A$ and $u,v\in W$, $u_{i},v_{i}\in W_{i}$ for $i=1,2$ be such that $[u,v]\cong[u_{1},v_{1}]\times[u_{2},v_{2}]$ (isomorphic as posets), and let $x\mapsto(x_{1},x_{2})$ be an isomorphism.
+Let $x,y\in[u,v]$ and $z,z^{\prime}$ be two amazing hypercube decompositions of $[u,v]$ which correspond to $(z_{1},z_{2})$ and $(z^{\prime}_{1},z^{\prime}_{2})$, respectively (see [5, Theorem 5.6]).
+Suppose $DS(z_{1},z_{1}^{\prime})=DS(z_{1}^{\prime},z_{1})$ and $DS(z_{2},z_{2}^{\prime})=DS(z_{2}^{\prime},z_{2})$.
+Then
+$DS(z,z^{\prime})=DS(z^{\prime},z).$
+Proof.
+By [5, Proposition 5.7]), for $p,p^{\prime},b,b^{\prime}\in[u,v]$, we have:
+  • $p\in W^{z}_{[u,v]}$ if and only if $p_{1}\in W^{z_{1}}_{[u_{1},v_{1}]}$ and $p_{2}\in W^{z_{2}}_{[u_{2},v_{2}]}$;
+  • $p^{\prime}\in W^{z^{\prime}}_{[u,v]}$ if and only if $p^{\prime}_{1}\in W^{z^{\prime}_{1}}_{[u_{1},v_{1}]}$ and $p^{\prime}_{2}\in W^{z^{\prime}_{2}}_{[u_{2},v_{2}]}$;
+  • $b\in W^{z^{\prime}\vee p}_{[p,v]}$ if and only if $b_{1}\in W^{z^{\prime}_{1}\vee p_{1}}_{[p_{1},v_{1}]}$ and $b_{2}\in W^{z^{\prime}_{2}\vee p_{2}}_{[p_{2},v_{2}]}$;
+  • $b^{\prime}\in W^{z^{\prime}\vee p^{\prime}}_{[p^{\prime},v]}$ if and only if $b^{\prime}_{1}\in W^{z_{1}\vee p^{\prime}_{1}}_{[p^{\prime}_{1},v_{1}]}$ and $b^{\prime}_{2}\in W^{z_{2}\vee p^{\prime}_{2}}_{[p^{\prime}_{2},v_{2}]}$.
+Hence, $DS(z_{1},z_{1}^{\prime})=DS(z_{1}^{\prime},z_{1})$ and $DS(z_{2},z_{2}^{\prime})=DS(z_{2}^{\prime},z_{2})$ imply the assertion. ∎
 Acknowledgments: We would like to thank Grant T.
-Barkley, Francesco Brenti, and Christian Gaetz for stimulating discussions we had during the Workshop “Bruhat order: recent developments and open problems”held at the University of Bologna from 15 April 2024 to 19 April 2024.
+Barkley, Francesco Brenti, Matthew Dyer, and Christian Gaetz for stimulating discussions we had during the Workshop “Bruhat order: recent developments and open problems”, held at the University of Bologna from 15 April 2024 to 19 April 2024.
 The authors are members of the I.N.D.A.M.
 group GNSAGA.
+Appendix: double hypercubes
+By Grant T. Barkley and Christian Gaetz
+In this appendix, we will prove an analog of Conjecture 3.6 for co-elementary intervals.
+We say $[u,v]\subset S_{n}$ is co-simple if the set of roots corresponding to the lower cover relations of $v$ in $[u,v]$ is linearly independent and co-elementary if $[u,v]$ is isomorphic as a poset to $[u^{\prime},v^{\prime}]$, where $[u^{\prime},v^{\prime}]$ is a co-simple interval in some symmetric group $S_{m}$ (see [1], where the poset dual notions of simple and elementary intervals are used).
+Every interval of the form $[u,w_{0}]$ is co-simple.
+In this appendix, rather than directly examining double shortcuts, we will study an analog that we call double hypercubes. Given an upper hypercube decomposition $z\in[u,v]$, we let
+[$\mathscr{H}^{z}_{[u,v]}$ :=$:=\{(\mathcal{H},p):p\in[z,v]\text{ and }\mathcal{H}\text{ is a hypercube in $[u,v]$ containing $u$,}$
+[…]
+We conjecture that for amazing hypercube decompositions, the map sending $(\mathcal{H},p)\in\mathscr{H}^{z}_{[u,v]}$ to $p$ is a bijection from [$\mathscr{H}^{z}_{[u,v]}$ to $W^{z}_{[u,v]}$.
+We will state our result in terms of [$\mathscr{H}^{z}_{[u,v]}$; together with the conjectured bijection, this would also imply cases of Conjecture 3.6.
+Definition 3.11.
+Let $z$ and $z^{\prime}$ be upper amazing hypercube decompositions of $[u,v]$. We denote by $DH(z,z^{\prime})$ the multiset
+$\{(|\mathcal{H}_{1}|+|\mathcal{H}_{2}|,b):(\mathcal{H}_{1},p)\in\mathscr{H}_{[u,v]}^{z}\text{ and }(\mathcal{H}_{2},b)\in\mathscr{H}^{z^{\prime}\vee p}_{[p,v]}\},$
+where $|\mathcal{H}|$ denotes the rank of the hypercube $\mathcal{H}$. An element of $DH(z,z^{\prime})$ is called a double hypercube.
+We now present Theorem 3.12, the main result of this appendix.
+This proves a direct analog of Conjecture 3.6, and gives a new proof of combinatorial invariance of $R$- and $\tilde{R}$-polynomials for co-elementary intervals via the analog of Theorem 3.9.
+Theorem 3.12.
+Let $[u,v]$ be a co-elementary interval.
+Let $z$ and $z^{\prime}$ be upper amazing hypercube decompositions which are strong in the sense of [1, Def.
+3.4].
+Then $DH(z,z^{\prime})=DH(z^{\prime},z)$ as multisets.
+Proof.
+We first note that the claim reduces immediately to the case of co-simple intervals, since everything in sight is preserved by poset isomorphism.
+We refer to [2, Chapter 5] or [1] for the notions of reflections labeling Bruhat graph edges, reflection orders, and increasing paths in the Bruhat graph with respect to a reflection order.
+We write $\mathcal{C}_{[x,y]}$ for the set of reflections labeling the coatoms of an interval $[x,y]$.
+We first prove the following lemma.
+Lemma 3.13.
+Let $[u,v]$ be co-simple and $z\in[u,v]$ be so that $[z,v]$ is diamond complete in $[u,v]$.
+Let $\leq_{1}$ and $\leq_{2}$ be reflection orders so that if $t\in\mathcal{C}_{[u,v]}\setminus\mathcal{C}_{[z,v]}$ and $t^{\prime}\in\mathcal{C}_{[z,v]}$, then $t\leq_{1}t^{\prime}$ and $t\leq_{2}t^{\prime}$.
+Then for any $p\in[z,v]$ and $k\in\mathbb{N}$, the number of $\leq_{i}$-increasing length-$k$ paths $\Gamma$ from $u$ to $p$ such that $\operatorname{supp}(\Gamma)\cap[z,v]=\{p\}$ is independent of $i\in\{1,2\}$.
+Proof of Lemma 3.13.
+For $p\in[z,v]$ write $a_{i}^{p,k}$ for the number of length-$k$ paths $\Gamma$ from $u$ to $p$ such that $\operatorname{supp}(\Gamma)\cap[z,v]=\{p\}$ which are increasing with respect to $\leq_{i}$.
+We induct on the size $|[u,p]|$.
+If $p=z$, then $a_{i}^{p,k}$ is the coefficient of $q^{k}$ in $\widetilde{R}_{u,p}$ (see Theorem 2.1), which is independent of $i$.
+For arbitrary $p\in[z,v]$, we have the identity
+$\sum_{\begin{subarray}{c}x\in[z,p]\\ k\in\mathbb{N}\end{subarray}}a_{i}^{x,k}q^{k}\widetilde{R}_{x,p}=\widetilde{R}_{u,p}.$
+The left-hand side counts pairs consisting of $\leq_{i}$-increasing paths from $u$ to $x$ which intersect $[z,p]$ at $\{x\}$ and $\leq_{i}$-increasing paths from $x$ to $p$.
+The right-hand side counts $\leq_{i}$ increasing paths from $u$ to $p$.
+The identity (Proof of Lemma 3.13.) then follows by the hypothesis on $\leq_{i}$, which implies that the concatenation of the two paths in a pair counted by the left-hand side is an increasing path (cf.
+the proof of [1, Corollary 4.6]).
+Now inductively we have that
+$\sum_{\begin{subarray}{c}x\in[z,p]\\ x\neq p\\ k\in\mathbb{N}\end{subarray}}a_{1}^{x,k}q^{k}\widetilde{R}_{x,p}=\sum_{\begin{subarray}{c}x\in[z,p]\\ x\neq p\\ k\in\mathbb{N}\end{subarray}}a_{2}^{x,k}q^{k}\widetilde{R}_{x,p}.$
+Evidently (Proof of Lemma 3.13.) implies
+$\sum_{\begin{subarray}{c}x\in[z,p]\\ k\in\mathbb{N}\end{subarray}}a_{1}^{x,k}q^{k}\widetilde{R}_{x,p}=\sum_{\begin{subarray}{c}x\in[z,p]\\ k\in\mathbb{N}\end{subarray}}a_{2}^{x,k}q^{k}\widetilde{R}_{x,p}.$
+Subtracting the two equations gives
+$\sum_{k\in\mathbb{N}}a_{1}^{p,k}q^{k}=\sum_{k\in\mathbb{N}}a_{2}^{p,k}q^{k},$
+from which the claim follows.
+We now pick two amazing hypercube decompositions $z,z^{\prime}\in[u,v]$.
+The remainder of the proof is simply an application of results from [1] to construct two reflection orders so that their increasing paths correspond with elements of $DH(z,z^{\prime})$ and $DH(z^{\prime},z)$, respectively.
+We then apply Lemma 3.13 to deduce the equality $DH(z,z^{\prime})=DH(z^{\prime},z)$.
+Note that [1] uses lower hypercube decompositions rather than upper, so the statements there are dual to the ones being used below.
+By [1, Lemma 4.5] and the proof of [1, Corollary 4.6], there is a reflection order $\leq_{z}$ with the property that any edge in $[z,v]\setminus[z^{\prime},v]$ precedes any edge in $[z,v]\cap[z^{\prime},v]$, and any edge in $[u,v]\setminus[z,v]$ precedes any edge in $[z,v]$.
+There is also an analogous reflection order $\leq_{z^{\prime}}$.
+Hence $\leq_{z}$ satisfies property (E) from [1] both for the hypercube decomposition $z\in[u,v]$ and for the hypercube decomposition $z\vee z^{\prime}\in[z,v]$.
+We deduce from two applications of [1, Theorem 3.10] that the $\leq_{z}$-increasing paths $\Gamma$ from $u$ to $v$ are in bijection with tuples $(\mathcal{H}_{1},p,\mathcal{H}_{2},b,\gamma)$, where $(\mathcal{H}_{1},p)$ is in $\mathscr{H}_{[u,v]}^{z}$ and $(\mathcal{H}_{2},b)$ is in $\mathscr{H}_{[p,v]}^{z\vee z^{\prime}}$ and $\gamma$ is a $\leq_{z}$-increasing path from $b$ to $v$.
+Furthermore the length of $\Gamma$ is $|\mathcal{H}_{1}|+|\mathcal{H}_{2}|+|\gamma|$, and from the proof of [1, Theorem 3.10], the path $\gamma$ is a subpath of $\Gamma$.
+Hence we deduce that the multiset
+$\{(|\Gamma|,b):\text{$\Gamma$ is $\leq_{z}$-increasing from $u$ to $b$},\leavevmode\nobreak\ \mathrm{supp}(\Gamma)\cap[z\vee z^{\prime},v]=\{b\}\}$
+coincides with the multiset $DH(z,z^{\prime})$. Similarly, the multiset
+$\{(|\Gamma|,b):\text{$\Gamma$ is $\leq_{z^{\prime}}$-increasing from $u$ to $b$},\leavevmode\nobreak\ \mathrm{supp}(\Gamma)\cap[z\vee z^{\prime},v]=\{b\}\}$
+coincides with the multiset $DH(z^{\prime},z)$. By Lemma 3.13, we deduce $DH(z,z^{\prime})=DH(z^{\prime},z)$.